OpenStudy (anonymous):

solve for x: 3√x-2 -8=8 |x-3|<7 4e^2x=12

OpenStudy (anonymous):

3√x-2 -8=8 3√x-2=8+8 x-2=16^3 x=4098

OpenStudy (anonymous):

|x-3|<7 x-3<7,x-3<-7 x<7+3 , x<3-7 x<10 , x<-4

OpenStudy (anonymous):

use a calculator for this 4e^2x=12 e^2x=7.389056099 4e^2x=12 x=0.4

OpenStudy (anonymous):

you're amazing :3 thanks